# Parsimonious Factor Models for Asymmetric Dependence in Multivariate Extremes

### v1

@@ 2.2 Extremes as Peaks-over-Thresholds
Another approach to modelling multivariate extremes characterises extreme events through threshold exceedances rather than through componentwise maxima. More precisely, assume the convergence in distribution of $\bm{a}_{n}^{-1}(\bm{M}_{n}-\bm{b}_{n})$ to $\bm{Y}$ as introduced above. Using the notation $\bm{X}\nleq\bm{b}_{n}$ to indicate that $\bm{X}\leq\bm{b}_{n}$ does not hold — i.e., there exists at least one $j\in\{1,\ldots,d\}$ such that $X_{j}>b_{n,j}$ — we have
@@
for $\bm{x}>\bm{\eta}$; otherwise $H(\bm{x})=0$ whenever $x_{i}<\eta_{i}$ for at least one $i\in\{1,\ldots,d\}$ (RootzenTajvidi2006). Here, $\wedge$ denotes the componentwise minimum operator. Using the representation (4), we obtain
@@
where $\tilde{\bm{\sigma}}=\bm{\sigma}-\bm{\xi}\bm{\mu}$, $\bm{\pi}=\bm{\tau}/\ell(\bm{\tau})$, and $\tau_{j}=\left(1+\xi_{j}\mu_{j}/\sigma_{j}\right)^{-1/\xi_{j}},j\in\{1,\ldots,d\}$ (Prop. 2 Rootzen.etal:2018b). When $\bm{\xi}=\bm{0}$, this representation simplifies to
@@
where $\partial\ell/\partial x_{k}$ denotes the first-order derivative of $\ell$ with respect to the $k$-th component and $\bm{e}_{j}$ is the $j$-th $d$-dimensional canonical vector (see Rootzen.etal:2018b, Proposition 17).
@@
$G_{j}(x)=\exp\left\{-\exp\{-x\}\right\},\quad x\in\mathbb{R},\quad j\in\{1,\ldots,d\}$
@@
[…]
@@ 3 A broad class of additive factor models for modelling extremes
$X_{i}=Z_{i}+(1/\alpha_{i})\cdot V_{0}\cdot\mathbf{1}(Z_{i}^{*}>c_{i}),\quad i\in\mathcal{I}=\{1,\ldots,d\},$ (6)
@@ 3.1 Extreme-value limiting distributions
Consider the random vector $\mathbf{X}=(X_{1},\ldots,X_{d})^{\top}$ defined in (6). Assume $V_{0}$ follows a unit exponential distribution, $c_{1}=c_{2}=\cdots=c_{d}=c_{0}$, $Z^{*}_{1}=Z^{*}_{2}=\cdots=Z_{d}^{*}=Z_{0}^{*}$, and $\rho_{k,0}=\mathrm{Cor}(Z_{k},Z_{0}^{*})$, $k\in\{1,\ldots,d\}$. Then the stable tail dependence function associated with $\mathbf{X}$ is given by
@@
$h(\bm{x})$ $=\phi_{d-1}^{\textrm{ESN}}\left[\left\{\frac{\lambda_{ik}}{2}+\frac{1}{\lambda_{ik}}\log\frac{x_{i}\Phi(\alpha_{i}\rho_{i,0}-c_{0})}{x_{k}\Phi(\alpha_{k}\rho_{k,0}-c_{0})}\right\}_{i\in\mathcal{I}_{-k}};\bar{\Sigma}^{(k)},\frac{\bar{\Sigma}^{(k)-1}\bm{\delta}_{k}}{\sqrt{1-\bm{\delta}^{\top}_{k}\bar{\Sigma}^{(k)-1}\bm{\delta}_{k}}},\frac{\alpha_{k}\rho_{k,0}-c_{0}}{\sqrt{1-\bm{\delta}^{\top}_{k}\bar{\Sigma}^{(k)-1}\bm{\delta}_{k}}}\right]$
$\quad\times\left\{\ell(\bm{1})x_{k}^{2}\prod_{i\in\mathcal{I}_{-k}}x_{i}\lambda_{ik}\right\}^{-1}.$ (8)
The extremal coefficient and the Pickands dependence function are also determined by the stable tail dependence function, via $\vartheta=\ell(\bm{1})$ and $A(\bm{w})=\ell\left(\bm{x}/\|\bm{x}\|_{1}\right)$.
@@ 4.1 Inference for Block Maxima data
$c_{\textrm{EV}}^{(i,j)}(u_{i},u_{j})$ $=(u_{i}u_{j})^{-1}\exp\left\{-\ell_{X_{i},X_{j}}(x_{i},x_{j})\right\}\cdot\left\{\frac{\partial\ell_{X_{i},X_{j}}(x_{i},x_{j})}{\partial x_{i}}\frac{\partial\ell_{X_{i},X_{j}}(x_{i},x_{j})}{\partial x_{j}}-\frac{\partial^{2}\ell_{X_{i},X_{j}}(x_{i},x_{j})}{\partial x_{i}\partial x_{j}}\right\}\,.$
@@ 5.1 Summer temperature maxima in Melbourne, Australia
    Submodel 1 from Section 3.2, with $c_{i}=0$ for $i\in\{1,\ldots,d\}$, corresponding to the Hüsler-Reiss distribution.
@@
    Submodel 1 from Section 3.2, with $c_{i}=\beta_{0}+\beta_{1}\bm{s}_{i1}+\beta_{2}\bm{s}_{i2}$, $i\in\{1,\ldots,d\}$, where $\{(\bm{s}_{i1},\bm{s}_{i2})\}_{i=1}^{d}$ denote the spatial coordinates of the $d=90$ locations.
@@
    Submodel 2 from Section 3.2, with $c_{i}=\beta_{0}+\beta_{1}\bm{s}_{i1}+\beta_{2}\bm{s}_{i2}$, $i\in\{1,\ldots,d\}$, where $\{(\bm{s}_{i1},\bm{s}_{i2})\}_{i=1}^{d}$ are the coordinates of the $d=90$ locations.
@@ A.1 Proof of Proposition 1
$\Pr(\mathbf{X}<\bm{w})=$ $\sum_{\mathcal{A}\in\{1,\ldots,d\},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}},\mathbf{Z}_{\mathcal{B}}<\bm{w}_{\mathcal{B}},\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}\}$
$=$ $\sum_{\mathcal{A}\in\{1,\ldots,d\},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}\}+o(1/n).$
@@
$\Pr(\mathbf{X}<\bm{w})$ $=\sum_{\mathcal{A}\in\{1,\ldots,d\},\mathcal{B}=\mathcal{A}^{c}}F_{0}\{\min_{i\in\mathcal{A}}(w_{i}\alpha_{i})\}\cdot\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}})+o(1/n)$
@@
$\ell_{\mathbf{X}}(\mathbf{x})=\lim_{n\to\infty}n\{1-\Pr(\mathbf{X}<\bm{w})\}=\sum_{\mathcal{A}\in\{1,\ldots,d\},\mathcal{B}=\mathcal{A}^{c}}\left(\max_{i\in\mathcal{A}}\frac{x_{i}}{\zeta_{i}}\right)\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}).$
@@ A.2 Proof of Proposition 2
$\Pr(\mathbf{X}<\bm{w})=$ $\sum_{\mathcal{A}\in\{1,\ldots,d\},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}},\mathbf{Z}_{\mathcal{B}}<\bm{w}_{\mathcal{B}},\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}\}$
$=$ $\sum_{\mathcal{A}\in\{1,\ldots,d\},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\tilde{\mathbf{X}}_{\mathcal{A}}=\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}}\}\cdot\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}})+o(1/n),$
@@
$n\{1-\Pr(\mathbf{X}<\bm{w})\}=\sum_{\mathcal{A}\in\{1,\ldots,d\},\mathcal{B}=\mathcal{A}^{c}}n\{1-\Pr(\tilde{\mathbf{X}}_{\mathcal{A}}<\bm{w}_{\mathcal{A}})\}\cdot\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}})+o(1).$

### v2

@@ 2.2 Extremes as Peaks-over-Thresholds
Another approach to modelling multivariate extremes characterises extreme events through threshold exceedances rather than through componentwise maxima. More precisely, assume the convergence in distribution of $\bm{a}_{n}^{-1}(\bm{M}_{n}-\bm{b}_{n})$ to $\bm{Y}$ as introduced above. Using the notation $\bm{X}\nleq\bm{b}_{n}$ to indicate that $\bm{X}\leq\bm{b}_{n}$ does not hold — i.e., there exists at least one $j\in\mathcal{I}=\{1,\ldots,d\}$ such that $X_{j}>b_{n,j}$ — we have
@@
for $\bm{x}>\bm{\eta}$; otherwise $H(\bm{x})=0$ whenever $x_{i}<\eta_{i}$ for at least one $i\in\mathcal{I}$ (RootzenTajvidi2006). Here, $\wedge$ denotes the componentwise minimum operator. Using the representation (4), we obtain
@@
where $\tilde{\bm{\sigma}}=\bm{\sigma}-\bm{\xi}\bm{\mu}$, $\bm{\pi}=\bm{\tau}/\ell(\bm{\tau})$, and $\tau_{j}=\left(1+\xi_{j}\mu_{j}/\sigma_{j}\right)^{-1/\xi_{j}},j\in\mathcal{I}$ (Prop. 2 Rootzen.etal:2018b). When $\bm{\xi}=\bm{0}$, this representation simplifies to
@@
[…]
@@
$G_{j}(x)=\exp\left\{-\exp\{-x\}\right\},\quad x\in\mathbb{R},\quad j\in\mathcal{I}$
@@
where $\Phi_{d}(\cdot;\Sigma)$ denotes the cumulative distribution function (cdf) of a $d$-dimensional normal distribution with covariance (or correlation) matrix $\Sigma$ and $\bar{\Sigma}^{(j)}$ is the partial correlation matrix with entries $(\bar{\Sigma}^{(j)})_{i,k}=\left(\lambda_{k,j}^{2}+\lambda_{i,j}^{2}-\lambda_{k,i}^{2}\right)/\left(2\lambda_{k,j}\lambda_{i,j}\right),\ i\neq k$. The stable tail dependence function is then
@@ 3 A broad class of additive factor models for modelling extremes
$X_{i}=Z_{i}+(1/\alpha_{i})\cdot V_{0}\cdot\mathbf{1}(Z_{i}^{*}>c_{i}),\quad i\in\mathcal{I},$ (6)
@@ 3.1 Extreme-value limiting distributions
Consider the random vector $\mathbf{X}=(X_{1},\ldots,X_{d})^{\top}$ defined in (6). Assume $V_{0}$ follows a unit exponential distribution, $c_{1}=c_{2}=\cdots=c_{d}=c_{0}$, $Z^{*}_{1}=Z^{*}_{2}=\cdots=Z_{d}^{*}=Z_{0}^{*}$, and $\rho_{k,0}=\mathrm{Cor}(Z_{k},Z_{0}^{*})$, $k\in\mathcal{I}$. Then the stable tail dependence function associated with $\mathbf{X}$ is given by
@@
$h(\bm{x})$ $=\phi_{d-1}^{\textrm{ESN}}\left[\left\{\frac{\lambda_{ik}}{2}+\frac{1}{\lambda_{ik}}\log\frac{x_{i}\Phi(\alpha_{i}\rho_{i,0}-c_{0})}{x_{k}\Phi(\alpha_{k}\rho_{k,0}-c_{0})}\right\}_{i\in\mathcal{I}_{-k}};\bar{\Sigma}^{(k)},\frac{\bar{\Sigma}^{(k)-1}\bm{\delta}_{k}}{\mathcal{S}_{k}},\frac{\alpha_{k}\rho_{k,0}-c_{0}}{\mathcal{S}_{k}}\right]$
$\quad\times\left\{\ell(\bm{1})x_{k}^{2}\prod_{i\in\mathcal{I}_{-k}}x_{i}\lambda_{ik}\right\}^{-1},$ (8)
where $\mathcal{S}_{k}=\sqrt{1-\bm{\delta}^{\top}_{k}\bar{\Sigma}^{(k)-1}\bm{\delta}_{k}}$. The extremal coefficient and the Pickands dependence function are also determined by the stable tail dependence function, via $\vartheta=\ell(\bm{1})$ and $A(\bm{w})=\ell\left(\bm{x}/\|\bm{x}\|_{1}\right)$.
@@ 4.1 Inference for Block Maxima data
$c_{\textrm{EV}}^{(i,j)}(u_{i},u_{j})$ $=\frac{\exp\left\{-\ell_{X_{i},X_{j}}(x_{i},x_{j})\right\}}{u_{i}u_{j}}\cdot\left\{\frac{\partial\ell_{X_{i},X_{j}}(x_{i},x_{j})}{\partial x_{i}}\frac{\partial\ell_{X_{i},X_{j}}(x_{i},x_{j})}{\partial x_{j}}-\frac{\partial^{2}\ell_{X_{i},X_{j}}(x_{i},x_{j})}{\partial x_{i}\partial x_{j}}\right\}\,.$
@@ 5.1 Summer temperature maxima in Melbourne, Australia
    Submodel 1 from Section 3.2, with $c_{i}=0$ for $i\in\mathcal{I}$, corresponding to the Hüsler-Reiss distribution.
@@
    Submodel 1 from Section 3.2, with $c_{i}=\beta_{0}+\beta_{1}\bm{s}_{i1}+\beta_{2}\bm{s}_{i2}$, $i\in\mathcal{I}$, where $\{(\bm{s}_{i1},\bm{s}_{i2})\}_{i=1}^{d}$ denote the spatial coordinates of the $d=90$ locations.
@@
    Submodel 2 from Section 3.2, with $c_{i}=\beta_{0}+\beta_{1}\bm{s}_{i1}+\beta_{2}\bm{s}_{i2}$, $i\in\mathcal{I}$, where $\{(\bm{s}_{i1},\bm{s}_{i2})\}_{i=1}^{d}$ are the coordinates of the $d=90$ locations.
@@ A.1 Proof of Proposition 1
$\Pr(\mathbf{X}<\bm{w})=$ $\sum_{\mathcal{A}\in\mathcal{I},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}},\mathbf{Z}_{\mathcal{B}}<\bm{w}_{\mathcal{B}},\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}\}$
$=$ $\sum_{\mathcal{A}\in\mathcal{I},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}\}+o(1/n).$
@@
$\Pr(\mathbf{X}<\bm{w})$ $=\sum_{\mathcal{A}\in\mathcal{I},\mathcal{B}=\mathcal{A}^{c}}F_{0}\{\min_{i\in\mathcal{A}}(w_{i}\alpha_{i})\}\cdot\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}})+o(1/n)$
@@
$\ell_{\mathbf{X}}(\mathbf{x})=\lim_{n\to\infty}n\{1-\Pr(\mathbf{X}<\bm{w})\}=\sum_{\mathcal{A}\in\mathcal{I},\mathcal{B}=\mathcal{A}^{c}}\left(\max_{i\in\mathcal{A}}\frac{x_{i}}{\zeta_{i}}\right)\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}).$
@@ A.2 Proof of Proposition 2
$\Pr(\mathbf{X}<\bm{w})=$ $\sum_{\mathcal{A}\in\mathcal{I},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}},\mathbf{Z}_{\mathcal{B}}<\bm{w}_{\mathcal{B}},\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}}\}$
$=$ $\sum_{\mathcal{A}\in\mathcal{I},\mathcal{B}=\mathcal{A}^{c}}\Pr\{\tilde{\mathbf{X}}_{\mathcal{A}}=\mathbf{Z}_{\mathcal{A}}+(1/\bm{\alpha}_{\mathcal{A}})\cdot V_{0}<\bm{w}_{\mathcal{A}}\}\cdot\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}})+o(1/n),$
@@
$n\{1-\Pr(\mathbf{X}<\bm{w})\}=\sum_{\mathcal{A}\in\mathcal{I},\mathcal{B}=\mathcal{A}^{c}}n\{1-\Pr(\tilde{\mathbf{X}}_{\mathcal{A}}<\bm{w}_{\mathcal{A}})\}\cdot\Pr(\mathbf{Z}^{*}_{\mathcal{A}}>\mathbf{c}_{\mathcal{A}},\mathbf{Z}^{*}_{\mathcal{B}}<\mathbf{c}_{\mathcal{B}})+o(1).$